# Decoding Noise in Nanofluidic Systems: Adsorption versus Diffusion Signatures in Power Spectra

### v1

@@ III.3 Full Case
We now consider the full problem where both diffusion and adsorption processes lead to fluctuations. Notably, in this case the results from the previous two sections do not add linearly. Instead, we observe a much more complicated picture, both in the analytical expressions in Eq. (12) and in the shapes and behavior of the PSDs, although some distinguishable “common” scaling regimes are reached. To build intuition, we therefore consider the following limiting cases:
@@ III.3.1 Diffusion Slower than Adsorption/Desorption
We first explore the case where diffusion is slower than both adsorption and desorption processes. Physically, slow diffusive timescales (relative to reaction kinetics) often appear in microfluidic systems. For example, the kinetics of antigen/antibody binding is fast compared to the time for antigens to diffuse across an immunoassay Hu, Gao, and Li (2007). Similarly, the adsorption/desorption of redox-active species in nanofluidic electronic devices can be fast with respect to their diffusion Zevenbergen et al. (2009).
@@
Results (both theory and simulations) are shown in Fig. 5, where the corresponding curves for a pure diffusion (dotted black line) or adsorption (dash-dotted black line) process are shown for comparison. The total PSD is close to the purely diffusive case, with a corner frequency around the diffusion timescale and a high frequency scaling of $-3/2$. It is perhaps counter-intuitive that the total PSD, $S_{N}$, does not show the corner frequency linked to adsorption given that the pure adsorption curve has a higher magnitude than the diffusion curve at high frequencies. This is a result of cross-correlations, which may be negative, i.e., even though $N=N_{\rm on}+N_{\rm off}$, the power spectra are not additive, $S_{N}\neq S_{\rm on}+S_{\rm off}$.
@@
Bound and free PSDs are similar to each other and both display two corner frequencies, one around the diffusive timescales and one at the binding timescales. With respect to scalings, the bound PSD appears to follow the pure adsorption case at high frequencies, while the free PSD deviates from this and seems to have a slightly less steep slope than $-2$ for very high frequencies, possibly even approaching the $-3/2$ regime. At low frequencies, both free and bound PSDs plateau at values that are below that of the purely diffusive case, again highlighting non-additivity. This is likely a scaling effect associated with the average number of bound/free particles being less than that of the particles in the channel. In the intermediate regime between the two corners, there is a transient scaling regime; this appears as $-3/2$ due to diffusion for the free PSD, but is sufficiently short that this may not always be obvious in experimental systems.
@@
The bound PSD is very similar to the spectrum of fluctuations in the current measured in Ref. [Zevenbergen et al., 2009]. In that nanofluidic system, it is likely the relevant quantity being probed by the current is the number of species adsorbed onto the electrodes, which means our results for the bound PSD are in good agreement. Our model rationalizes these results in the context of other fluctuating quantities in the system. Finally, we note that the ordering of adsorption and desorption timescales does not significantly affect the shape of the power spectra; the complementary case where adsorption is faster than desorption ($k_{\rm on}\gg k_{\rm off}$) is discussed in Appendix D, in particular in Fig. 8.
@@ III.3.2 Diffusion Faster than Adsorption/Desorption
We now turn to the situation where diffusion is faster than adsorption/desorption. Physically, this ordering of timescales is common in nanopores, both biological and engineered. For example, it occurs naturally in the transport of sugars through maltoporins (LamB transmembrane proteins) Kullman, Winterhalter, and Bezrukov (2002); Nekolla, Andersen, and Benz (1994) and also in the use of adapters, such as beta cyclodextrin, in nanopore sensors for detection of organic molecules Gu, Cheley, and Bayley (2001). Nanoscale electrical devices where analytes reversibly bind to the electrodes may also involve fast diffusion relative to the adsorption timescales Kätelhön et al. (2014).
@@
Representative spectra are shown in Fig. 6. Here, we only consider the case where $k_{\rm on}=k_{\rm off}$, because varying $k_{\rm on}$ and $k_{\rm off}$ does not significantly affect the shape or behavior of the PSDs, provided both correspond to timescales much shorter than that for diffusion.
Figure 6: Fast diffusion, slow adsorption/desorption: Case 2 – $\tau_{\rm cha}\ll\tau_{\rm ads}\sim\tau_{\rm des}$ Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$, $k_{\rm on}=k_{\rm off}=0.008/\tau_{0}$. The color/line code is the same as that of Fig. 5.
@@
We observe that the total PSD is more sensitive to the adsorption/desorption process compared to the previous case. In particular, it now shows two corners and two slopes, with the low-frequency corner determined by the adsorption/desorption timescale and high frequency corner by the diffusive timescale(s). Again, there is a transient intermediate scaling regime between the two corners. The slopes of all three PSDs at high frequencies seem to take the same values as in the pure adsorption or pure diffusion cases respectively (i.e. $-2$ and $-3/2$), depending on the PSD.
@@
In this case, the total and free PSDs have similar shapes whereas previously the free and bound PSDs were similar. Here, with fast diffusion, the number of bound particles is less dependent on the diffusive timescale as a rate-limiting step. Instead, at high frequencies and short timescales, the number of bound particles remains roughly constant, and the total number of particles reflects changes in the number of free particles. This results in a $-3/2$ scaling in total and free PSDs due to particles diffusing in and out of the reservoirs. At low frequencies, the free and total PSDs show corners at the adsorption/desorption timescale, since at longer times particles can equilibrate across the whole system (diffusing, adsorbing, and unadsorbing) and so the fluctuations become uncorrelated.
@@
In the context of Ref. [Zevenbergen et al., 2009], their observation of $-3/2$ scaling at intermediate frequencies and $-2$ at high frequencies does not appear in this case in any of the measured fluctuating quantities. As such, we can conclude that the diffusive timescales must be slower than the adsorption/desorption timescales for their system. This highlights that the ordering of slopes in the PSD can be indicative of the relative timescales in the system, a relatively simple criterion to observe.
@@ III.3.3 Further Cases
Figure 7: Left: physical systems corresponding to different diffusion/adsorption regimes. Right: the corresponding shapes of the PSDs identified in this work.
@@
As highlighted in Table 1, other experimental systems may exhibit further different timescale separations, where all timescales are comparable (Case 3) or where adsorption is much faster than all other processes (Case 4). In these scenarios, the spectra obtained are again non-linear combinations of the pure diffusion and pure adsorption cases, and can either be quite similar to either of these pure cases (Case 3) or exhibit entirely peculiar shapes (Case 4). More detailed investigations are reported in Appendix D.
@@
To summarize the different possibilities for the PSD shapes, we build a phase diagram – see Fig. 7, right-hand side – comparing the diffusive and the adsorption/desorption timescales respectively. A key takeaway is that there exists only a few special cases in which the PSDs exhibit two corners, related to two distinct slopes, while in all other cases, the PSDs appear to have an approximately Lorentzian, smooth, shape, with a single well-defined or somewhat smeared out scaling regime. These special “two distinct slopes” cases are only observed in case 1 of slow diffusion compared to binding timescales, for bound and free PSDs; and in case 2 of fast diffusion compared to binding timescales, for total and free PSDs. The free PSD thus always senses the two regimes, but the total and bound PSDs do not. Depending on what particles the experimental measurement most represents, there are thus only a few restricted cases where one could see these two corners. Finally, depending on the ratio between the different time scales, the scaling laws probed may either be $1/f^{3/2}$ or $1/f^{2}$ at high frequencies.
@@ IV Discussion and Conclusions
Spectra show a number of characteristic scalings that reflect key physical mechanisms and properties of the system, as outlined in Table 2. By considering the 1D case, we have also eliminated the effect of geometry on scaling Gravelle, Netz, and Bocquet (2019); Robin et al. (2023a) and significantly simplified interpretation of the results. Interestingly, our analytic results indicate that there are only a small number of robust scalings which emerge directly from physical mechanisms, with slopes of $-1/2$, $-3/2$, and $-2$ respectively. Observations of other scalings are the result of intermediate transitions between regimes or combination of two physical processes. The scaling of the power spectra thus also provides valuable insight into the dominant mechanism generating number fluctuations across different frequency ranges.
@@
plateau low $f$ finite system size, finite global decorrelation time
$1/f^{1/2}$ intermediate $f$ diffusive re-entrance effects
@@
$1/f^{2}$ high $f$ independent events, like adsorption/desorption
@@
Real systems are clearly more complex than the situation considered here, but we can find scenarios in experiments that correspond to the different separation of timescales between diffusion and adsorption that we have studied. These are summarized in Fig. 7, left-hand side, and clearly span the entire phase diagram, indicating that there is no general shape to be expected from a PSD even with such simple ingredients. The ordering of slopes can also be used in the first instance to determine relative timescales of diffusion and adsorption: a relatively simple criterion to check. An interesting question for all of these cases is which of our modeled PSDs – total, free, or bound – best reflects the physical quantity measured experimentally. For example, nanopore systems are often investigated via ionic current fluctuations. Here, the current is a measure of the ions in solution traveling through the pore; therefore it could be reasonable to use the total PSD as an analogue for fluctuations in the current, assuming a degree of correlation between the pore conductance and total particles blocking the channel (adapter, sugar, etc) Secchi et al. (2016). In assays or nano-electrical devices, by design, the adsorbed particles may have the most significant effect on a macroscopic fluctuating signal, in which case the bound PSD is most relevant.
@@ Acknowledgements.
The authors have to conflicts to disclose.
@@ B.0.1 Number of adsorbed particles
$\hat{C}_{\rm on}$ $(s)=\frac{2\mathcal{N}}{s}\Bigg{[}\frac{sk_{\rm on}A_{C}}{m_{1}(s+k_{\rm off})}$ (32)
$\Big{(}e^{L_{0}m_{1}/2}-e^{-L_{0}m_{1}/2}\Big{)}-\frac{q_{0}k_{\rm off}L_{0}}{2(s+k_{\rm off}+k_{\rm on})}\Bigg{]}.$

### v2

@@ III.3 Full Case
We now consider the full problem where both diffusion and adsorption processes lead to fluctuations. Notably, in this case the results from the previous two sections do not add linearly, i.e. we have $S_{N}(f)\neq S_{\rm on}(f)+S_{\rm off}(f)$. In practice, this is due to the presence of cross-correlations in the fluctuations, and the spectrum of the total number of particles in the channel is the sum of 4 terms
$S_{N}(f)=S_{\rm on}(f)+S_{\rm off}(f)+S_{\rm{on\rightarrow off}}(f)+S_{\rm{off\rightarrow on}}(f).$ (19)
The cross-correlation terms are in general non-trivial (see Appendix B). Although we now observe a much more complicated picture, both in the analytical expressions in Eq. (12) and in the shapes and behavior of the PSDs, some distinguishable “common” scaling regimes are reached which we discuss below. To build intuition, it is useful to consider the following limiting cases:
@@
Figure 6: Fast diffusion, slow adsorption/desorption: Case 2 – $\tau_{\rm cha}\ll\tau_{\rm ads}\sim\tau_{\rm des}$ Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$, $k_{\rm on}=k_{\rm off}=0.05/\tau_{0}$. The color/line code is the same as that of Fig. 5.
@@ III.3.1 Diffusion Slower than Adsorption/Desorption
We first explore the case where diffusion is slower than both adsorption and desorption processes. Physically, this could manifest as the adsorption/desorption of redox-active species in nanofluidic electronic devices, which can be fast with respect to their diffusion Zevenbergen et al. (2009).
@@
Results (both theory and simulations) are shown in Fig. 5, where the corresponding curves for a pure diffusion (dotted black line) or adsorption (dash-dotted black line) process are shown for comparison. The total PSD is close to the purely diffusive case, with a corner frequency around the diffusion timescale and a high frequency scaling of $f^{-3/2}$. It is perhaps counter-intuitive that the total PSD, $S_{N}$, does not show the corner frequency linked to adsorption given that the pure adsorption curve has a higher magnitude than the diffusion curve at high frequencies. Mathematically, this is a result of cross-correlations, which may be negative, i.e. the power spectra are not additive, $S_{N}\neq S_{\rm on}+S_{\rm off}$ (see Appendix B). Physically, the reason for this is that adsorption and desorption processes are so fast that they mostly affect the relative distribution of bound or free particles, but these processes have “averaged” out before a particle may escape the channel by diffusion and contribute fluctuations to the total PSD.
@@
Bound and free PSDs are similar to each other and both display two corner frequencies, one around the diffusive timescales and one at the binding timescales. At low frequencies, both PSDs plateau at values which sum to less than that of the purely diffusive case, again highlighting non-additivity. In the intermediate regime between the two corners, there is a transient scaling regime; this appears as $f^{-3/2}$ due to diffusion for the free PSD, but is sufficiently short that this may not always be obvious in experimental systems. At high frequencies, the bound PSD is dominated by a $f^{-2}$ scaling, which intuitively makes sense as the faster processes of adsorption and desorption dominate on short timescales.
There is an important distinction between the bound and free PSDs; at high frequencies, although the free PSD follows a $f^{-2}$ scaling just above the adsorption/desorption rates, it eventually returns to a $f^{-3/2}$ scaling. This is not obvious over a short range of frequencies, as the deflection is very slow, and occurs above $f\sim L_{0}^{2}k_{\rm on}^{2}/D\pi\sim\tau_{\rm cha}/\tau_{\rm ads}^{2}$. In all cases, in fact, the free PSD returns to a $f^{-3/2}$ scaling at high frequencies as a limiting case. This is due to the fact that if particles are “close” enough to the channel boundary, they will diffusively enter/exit the channel at infinitely fast timescales, and hence a diffusive $f^{-3/2}$ scaling has to be recovered at high enough frequencies. Crucially, the slow transition between these scaling regimes would not be clear in an experiment which can only probe a few decades of frequency space, so it is possible that the $f^{-3/2}$ scaling is missed entirely or mistaken for having an exponent between 1.5-1.5 and 2.0-2.0. The absence of a sharp corner also masks this underlying shift in scalings.
@@
The bound PSD is very similar to the spectrum of fluctuations in the current measured in Ref. [Zevenbergen et al., 2009]. In that nanofluidic system, it is likely that the relevant quantity being probed by the current is the number of species adsorbed onto the electrodes, which means our results for the bound PSD are in good agreement. Our model rationalizes these results in the context of other fluctuating quantities in the system. Finally, we note that the ordering of adsorption and desorption timescales does not significantly affect the shape of the power spectra; the complementary case where adsorption is faster than desorption ($k_{\rm on}\gg k_{\rm off}$) is discussed in Appendix D, in particular in Fig. 9.
@@ III.3.2 Diffusion Faster than Adsorption/Desorption
Figure 7: (Top) Characteristic shapes and scalings of the PSDs identified in this work. (Bottom) Corresponding physical systems in different diffusion/adsorption regimes.
@@
We now turn to the situation where diffusion is faster than adsorption/desorption. This case is the one considered in Ref. [Berezhkovskii, Pustovoit, and Bezrukov, 2002]. Physically, this ordering of timescales is common in nanopores, both biological and engineered. For example, it occurs naturally in the transport of sugars through maltoporins (LamB transmembrane proteins) Kullman, Winterhalter, and Bezrukov (2002); Nekolla, Andersen, and Benz (1994) and also in the use of adapters, such as beta cyclodextrin, in nanopore sensors for detection of organic molecules Gu, Cheley, and Bayley (2001). Nanoscale electrical devices where analytes reversibly bind to the electrodes may also involve fast diffusion relative to the adsorption timescales Kätelhön et al. (2014).
@@
Representative spectra are shown in Fig. 6. Here, we only consider the case where $k_{\rm on}=k_{\rm off}$, because varying $k_{\rm on}$ and $k_{\rm off}$ does not significantly affect the shape or behavior of the PSDs.
@@
We observe that the total PSD is more sensitive to the adsorption/desorption process compared to Case 1 where diffusion was slower than adsorption/desorption. In particular, it now shows two corners and two slopes, with the low-frequency corner determined by the adsorption/desorption timescale and high frequency corner by the diffusive timescale(s). Again, there is a transient intermediate scaling regime between the two corners.
@@
In this case, the total and free PSDs have similar shapes whereas previously the free and bound PSDs were similar. Here, with fast diffusion, the number of bound particles is less dependent on the diffusive timescale as a rate-limiting step, and hence the bound PSD only exhibits a single slope that is similar to the purely adsorbing case. At high frequencies and short timescales, the number of bound particles remains roughly constant, but the total number of particles and of free particles change due to diffusive entrance/exits effects. This results in a $f^{-3/2}$ scaling for both the free and total PSDs at high frequencies; the situation is similar to the purely diffusive case but with fewer diffusing particles, since some are bound. At low frequencies, the free and total PSDs show corners at the adsorption/desorption timescale, since at longer times particles can equilibrate across the whole system, diffusing, adsorbing, and desorbing.
@@
In the context of Ref. [Zevenbergen et al., 2009], their observation of $f^{-3/2}$ scaling at intermediate frequencies and $f^{-2}$ at high frequencies does not appear in this case in any of the measured fluctuating quantities. As such, we can conclude that the diffusive timescales must be slower than the adsorption/desorption timescales for their system. This highlights that the ordering of slopes in the PSD can be indicative of the relative timescales in the system, a relatively simple criterion to observe.
@@ III.3.3 Further Cases
As highlighted in Table 1, other experimental systems may exhibit further different timescale separations, where all timescales are comparable (Case 3) or where adsorption is much faster than all other processes (Case 4). In these scenarios, the spectra obtained are again non-linear combinations of the pure diffusion and pure adsorption cases, and can either be quite similar to either of the two pure cases (Case 3) or exhibit entirely peculiar shapes (Case 4). More detailed investigations are reported in Appendix D. Overall, the two corners are not observed in these further cases.
III.4 Key Takeaways
To summarize the different possibilities for the PSD shapes and scalings, we build a phase diagram – see Fig. 7 – comparing the diffusive and adsorption/desorption timescales.
@@ III.4 Key Takeaways
The first key takeaway concerns PSD shapes. In most cases, the PSD appears to have an approximately Lorentzian, smooth, shape with a single well-defined or somewhat smeared out scaling regime that is more difficult to interpret. Yet, there exist a few special cases in which the PSDs can exhibit two corners, related to two distinct slopes. These special “two corner” cases are only observed in some but not all of the spectra in the cases where there is a clear separation of timescales between the diffusive processes and adsorption/desorption processes (Cases 1 and 2). The PSD of free particles always exhibits the two corners in these cases, which makes sense since it is the only quantity for which both adsorption and diffusion affect number fluctuations directly. In contrast, the PSD of the total number of particles only exhibits two corners when diffusion is fast compared to binding, a condition by which it is indirectly able to sense the binding process. Reciprocally, the PSD of the number of bound particles only exhibits two corners when diffusion is slow compared to binding, a condition by which is is indirectly able to sense diffusion.
The second key takeaway concerns the characteristic scalings that one can observe in the PSDs. The total and free PSDs always tend to a $f^{-3/2}$ scaling at high frequencies because they sense in particular particles that are diffusing and hence diffusive boundary crossings at short timescales. In contrast, the bound PSD always tends to a $f^{-2}$ scaling at high frequencies because it only senses adsorption and desorption processes to first order. In the cases (1 and 2) where there is a clear separation of timescales between diffusive and adsorption/desorption processes, two distinct scalings emerge (on the relevant curves) separated by a corner. These scalings appear necessarily in the order corresponding to the timescale separation: e.g. if diffusion is the slower process, the $f^{-3/2}$ scaling will appear at intermediate frequencies, and $f^{-2}$ at high frequencies, and reciprocally if diffusion is the faster process. Naturally, because the free PSD has to decay to $f^{-3/2}$ at the highest frequencies but can exhibit a $f^{-2}$ scaling at high frequencies in the case of slow diffusion, the free PSD may exhibit two slopes at high frequencies separated by a slow transition, not a corner, which could mask the true scalings experimentally and result in effective fits of the slope between 1.5-1.5 to 2.0-2.0.
Finally, it is clear that there is significant non-additivity in these spectra and that, in general, there is not a straightforward route to construct the total PSD from the underlying spectra linked to pure diffusion and pure adsorption processes or from the bound and free PSDs. The physical reason for this is that processes are immediately coupled – as soon as a particle desorbs it can diffuse – and thus noise properties are convoluted.
More generally, we expect the phenomenology uncovered, in particular the presence of two corners or not according to the relative separation of timescales, should still hold in 3D or more complex geometries. To support this claim, we extend our simulations to a simple 3D system. As in Sec. III.1, we run Brownian dynamics simulations in a 3D cube with an adsorbing region at the center. Results are shown in Fig. 8. In this example, we choose the parameters to align with Case 1 of slow diffusion and fast adsorption/desorption. As in the 1D system, two slopes and two corner frequencies are observable in the bound and free PSDs due to the separation of timescales. Moreover, the slope below the corner corresponding to the adsorption/desorption timescales scales as $f^{-2}$, while at low frequencies below the diffusive timescales it is closer to $f^{-3/2}$. At high frequencies, the free PSD can be seen to tend to a $f^{-3/2}$ scaling. This is in agreement with our takeaways from the theoretical 1D case and demonstrates the robustness of our results to changes in geometry and dimension.
Figure 8: Simulated power spectral density of number fluctuations in the channel for a 3D system, for fluctuations in the total number (purple), the bound number (red), and the freely diffusing number (blue). We use $L=20\ell_{0}$ and $k_{\rm on}=k_{\rm off}=375/\tau_{0}$.
@@ IV Discussion and Conclusions
Spectra show a number of characteristic scalings that reflect key physical mechanisms and properties of the system, as outlined in Table 2. By considering the 1D case, we have also eliminated the effect of geometry on scaling Gravelle, Netz, and Bocquet (2019); Robin et al. (2023a) and significantly simplified interpretation of the results. Interestingly, our analytic results indicate that there are only a small number of robust scalings which emerge directly from diffusive and adsorption mechanisms, with slopes of $f^{-3/2}$ and $f^{-2}$ respectively. Observations of other scalings are the result of intermediate transitions between regimes. The scaling of the power spectra thus also provides valuable insight into the dominant mechanism generating number fluctuations across different frequency ranges.
@@
plateau low $f$ finite system size
$1/f^{1/2}$ intermediate $f$ 1D re-entrances effects
@@
$1/f^{2}$ high $f$ independent events like adsorption and desorption
@@
Real systems are clearly more complex than the situation considered here, but we can find scenarios in experiments that correspond to the different separations of timescales between diffusion and adsorption that we have studied. These are summarized in Fig. 7, left-hand side, and clearly span the entire phase diagram, indicating that there is no general shape to be expected from a PSD even with such simple ingredients. The ordering of slopes can also be used in the first instance to determine relative timescales of diffusion and adsorption: a fairly simple criterion to check. An interesting question for all of these cases is which of our modeled PSDs – total, free, or bound – best reflects the physical quantity measured experimentally. For example, nanopore systems are often investigated via ionic current fluctuations. Here, the current is a measure of the ions in solution traveling through the pore; therefore it could be reasonable to use the total PSD as an analogue for fluctuations in the current, assuming a degree of correlation between the pore conductance and total particles blocking the channel (adapter, sugar, etc) Secchi et al. (2016). In nano-electrical devices, by design, the adsorbed particles may have the most significant effect on a macroscopic fluctuating signal, in which case the bound PSD is most relevant.
@@ Acknowledgements.
The authors have no conflicts to disclose.
@@ B.0.1 Number of adsorbed particles
$\hat{C}_{\rm on}$ $(s)=2\mathcal{N}\Bigg[\frac{k_{\rm on}A_{C}}{m_{1}(s+k_{\rm off})}$ (33)
$\Big(e^{L_{0}m_{1}/2}-e^{-L_{0}m_{1}/2}\Big)-\frac{q_{0}k_{\rm off}L_{0}}{2s(s+k_{\rm off}+k_{\rm on})}\Bigg].$
@@ B.0.2 Number of unadsorbed particles
B.0.3 Cross-correlations
In general, we can also calculate correlation functions for the numbers of particles which switch from bound to free, $C_{\rm on\rightarrow off}(t)$, and vice versa, $C_{\rm off\rightarrow on}(t)$. These cross-correlations are non-trivial, which means that in most cases one cannot simply sum the contributions to the total PSD of free and bound number fluctuations, i.e. $S_{N}(f)\neq S_{\rm on}(f)+S_{\rm on}(f)$. The cross-correlations can be calculated using the same method as Sec. II.2:
$C_{\rm on\rightarrow off}(t)$ $=\langle N_{\rm off}(t)N_{\rm on}(0)\rangle$ (37)
$=\mathcal{N}\psi_{\rm on\rightarrow off}(t)$
where $\psi_{\rm on\rightarrow off}(t)$ represents the probability that a particle starting adsorbed in the channel is unadsorbed in the channel at time $t$. The initial condition is now that $p_{0}=0$. Again we take the time derivative and Laplace transform and using Eqs. (28) and (29), we obtain:
$\hat{C}_{\rm on\rightarrow off}$ $(s)=2\mathcal{N}\Bigg[A_{C}\Big(e^{L_{0}m_{1}/2}-e^{-L_{0}m_{1}/2}\Big)$ (38)
$\Big(\frac{m_{1}}{m_{0}^{2}}-\frac{k_{\rm on}}{m_{1}(s+k_{\rm off})}\Big)+\frac{q_{0}k_{\rm off}L_{0}}{2s(s+k_{\rm off}+k_{\rm on})}\Bigg].$
The other cross-correlations follows:
$C_{\rm off\rightarrow on}(t)$ $=\langle N_{\rm on}(t)N_{\rm off}(0)\rangle$ (39)
$=\mathcal{N}\psi_{\rm off\rightarrow on}(t)$
now with $q_{0}=0$. As above:
$\hat{C}_{\rm off\rightarrow on}$ $(s)=2\mathcal{N}\Bigg[A_{C}\Big(e^{L_{0}m_{1}/2}-e^{-L_{0}m_{1}/2}\Big)$ (40)
$\Big(\frac{k_{\rm on}}{m_{1}(s+k_{\rm off})}\Big)+\frac{p_{0}k_{\rm on}L_{0}}{2s(s+k_{\rm off}+k_{\rm on})}\Bigg].$